# Zero-Shot Transferable Solution Method for Parametric Optimal Control Problems

### v1

@@ Abstract
This paper presents a transferable solution method for optimal control problems with varying objectives using function encoder (FE) policies. Traditional optimization-based approaches must be re-solved whenever objectives change, resulting in prohibitive computational costs for applications requiring frequent evaluation and adaptation. The proposed method learns a reusable set of neural basis functions that spans the control policy space, enabling efficient zero-shot adaptation to new tasks through either projection from data or direct mapping from problem specifications. The key idea is an offline–online decomposition: basis functions are learned once during offline imitation learning, while online adaptation requires only lightweight coefficient estimation. Numerical experiments across diverse dynamics, dimensions, and cost structures show our method delivers near-optimal performance with minimal overhead when generalizing across tasks, enabling semi-global feedback policies suitable for real-time deployment.
@@ II Background
where $L:\mathbb{R}^{n}\times\mathbb{R}^{m}\times[0,T]\to\mathbb{R}$ is the running cost, $G:\mathbb{R}^{n}\to\mathbb{R}$ is the terminal cost, and $\mathcal{U}$ is the space of admissible control functions, which we assume to be sufficiently regular [16, Sec. I.3, I.8-9 ]. For simplicity and without loss of generality, we use $\boldsymbol{\eta}$} to denote the conditional variable that captures task-specific variations in $\mathcal{J}$, such as target locations, terrain types, or changes in control penalties. In practice, $\boldsymbol{\eta}$} can be complex, high-dimensional or implicit. Each choice of $\boldsymbol{\eta}$} defines a distinct optimal control problem.
@@
for each specific task. Having the control $\bf u$} to depend on both the current state ${\bf x}(t)$ and time $t$ allows for control evaluation for any state-time pair, making it more flexible and particularly useful in problems where the initial state ${\bf x}_{0}$ can vary. Classical results show that under suitable regularity assumptions, existence of optimal controls in feedback form can be guaranteed; see, e.g., [16, Sec. I.5], or [6].
@@
Function encoder (FE) [23] provides a principled framework for representing and transferring tasks in Hilbert spaces by learning a finite set of neural network basis functions. Given a Hilbert space $\mathcal{H}$, the FE learns basis functions $\{\phi_{1},\phi_{2},\dots,\phi_{p}\}$, parameterized by neural networks, such that any function $f\in\mathcal{H}$ can be approximated as
$f({\bf x})\approx\sum_{j=1}^{p}c_{j}\phi_{j}({\bf x};{\boldsymbol{\theta}}_{j}),$ (4)
@@
Online computation of coefficients for a given new problem $f$ is obtained by solving a least-squares (LS) problem
${\bf c}=\operatorname*{arg\,min}_{{\bf c}\in\mathbb{R}^{p}}\|f-\sum_{j=1}^{p}c_{j}\phi_{j}\|_{\mathcal{H}}^{2}.$ (5)
@@
${\bf c}=\begin{bmatrix}\langle\phi_{1},\phi_{1}\rangle_{\mathcal{H}}&\cdots&\langle\phi_{1},\phi_{p}\rangle_{\mathcal{H}}\\ \vdots&\ddots&\vdots\\ \langle\phi_{p},\phi_{1}\rangle_{\mathcal{H}}&\cdots&\langle\phi_{p},\phi_{p}\rangle_{\mathcal{H}}\end{bmatrix}^{-1}\begin{bmatrix}\langle f,\phi_{1}\rangle_{\mathcal{H}}\\ \vdots\\ \langle f,\phi_{p}\rangle_{\mathcal{H}}\end{bmatrix}$, (6)
@@
Let $K\subset\mathbb{R}^{n}$ be compact and let $\mathcal{H}=\{f:K\to\mathbb{R}^{n}\mid\|f\|_{\mathcal{H}}<\infty\}$ be a separable Hilbert space. For any continuous $f\in\mathcal{H}$ and any $\epsilon>0$, there exist neural network basis functions $\{\phi_{j}\}_{j=1}^{P}$ and coefficients $\{c_{j}\}_{j=1}^{P}$ such that
$\Big\|f-\sum_{j=1}^{P}c_{j}\,\phi_{j}\Big\|_{\mathcal{H}}<\epsilon\,\|f\|_{\mathcal{H}}.$
@@ 7 ${\boldsymbol{\gamma}}\leftarrow{\boldsymbol{\gamma}}-\beta\,\nabla_{{\boldsymbol{\gamma}}}L$;
We note that while Theorem 1 guarantees that the bias of the policy adaptation can be arbitrarily small, (8) is only a discrete formulation of (5), where the inner product in $\mathcal{H}$ is approximated by finite measurements. For a given task, denote the minimizers of (5) and (8) be $\bf c$} and $\bf c$$^{\star}$, respectively. They generally differ due to sampling size and measurement noise.
@@
Under standard assumptions for LS regression with fixed design. Denote by $\Phi\in\mathbb{R}^{M\times p}$ the matrix whose $i$-th row is $\phi({\bf x}_{i},t_{i})^{\top}$ and $B=\tfrac{1}{M}\Phi^{\top}\Phi$ the Gram matrix. Assume sub-Gaussian noise in measurements and $B$ full rank, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\|{\bf c}-{\bf c}^{\star}\|_{2}\;\leq\;\frac{\sigma}{\sqrt{\lambda_{\min}(B)}}\,\sqrt{\frac{p+\log(1/\delta)}{M}}\;=\;\mathcal{O}\!\big(M^{-1/2}\big),$
for some constant $\sigma$.
The proof follows [41, Thm. 2.2, 2.3], the details of which we omit here. This result, together with Theorem 1, suggests asymptotic convergence of unbiased policy prediction as the number of bases and samples grows, validating our proposed approach in principle for both offline training and online inference steps. Empirical validation on function encoder is also available at [23].
@@ IV RELATED WORK
ML amortizes solve time into an offline phase to produce feedback policies evaluable online—bridging local trajectory optimization and global HJB solvers. Notably, these learned policies are usually semi-global: by opting for sampling instead of meshing, they provide reliable feedback over broad regions of the state–time domain, making them scalable, though they remain task-specific in their standard form. Data-driven approaches such as [19, 24, 12, 13] train policies on solver-generated trajectories and can achieve fast convergence and high accuracy. Model-based and physics-informed approaches [18, 36, 43, 15, 35, 14, 26, 20, 10, 33] leverage OC structure (objective terms, HJB residuals, Koopman representation, etc.) to provide robustness and reduce data requirements in training, though the resulting optimization can be considerably more challenging. Model-free RL [44, 21] learns feedback without an explicit mathematical model but as a trade-off is often sample-inefficient and may underperform model-based approaches [45]. While ML-based approaches have seen extraordinary success in solving optimal control problems, repurposing a trained model to new tasks efficiently and accurately remains an open challenge. This work aims to address that gap.
@@ V Numerical Experiments
Our FE implementation closely follows [23]. We use multi-head, multi-layer perceptrons (MLPs) for basis function parameterization. Exact training details vary depending on the example. For data preparation and ground truth comparison we rely on standard discretization-based open-loop solvers using SciPy or CasADi [4]. Our code is implemented in PyTorch [38] and will be released upon publication. All experiments were conducted on an NVIDIA A100 GPU.
@@
Evaluation
True
Objective Loss
Predicted
Objective Loss
Inference
Method
@@ V-B Quadcopter Path Planning with Different Targets
Evaluation
True
Objective Loss
Predicted
Objective Loss
Inference
Method
@@ V-C Bicycle Control Under Different Obstacle Configurations
TABLE III: Numerical results for the obstacle generalization experiments using the bicycle model. We display the prediction and ground truth for the cumulative control and obstacle cost, as well as terminal state deviation measured in both $x$ and $y$ direction, all results are evaluated and averaged over multiple new problem setting unseen in training.
@@
We illustrate the numerical results in Table III. Note that the learned model transfers accurately to new problem settings using only limited data for inference. In particular, the controller can learn to avoid the obstacle for arbitrary placement and reach the target state with high precision, demonstrating the capability of our proposed approach. We present qualitative results in Figure 5. Here, we display the results for $3$ worst performing subproblems out of all tested cases. Notice that even in the worst performing cases, the learned model can provide reasonable and accurate guidance to the system, as demonstrated in the close match between the predicted and ground truth trajectories.

### v2

@@ Abstract
This paper presents a transferable solution method for optimal control problems with varying objectives using function encoder (FE) policies. Traditional optimization-based approaches must be re-solved whenever objectives change, resulting in prohibitive computational costs for applications requiring frequent evaluation and adaptation. The proposed method learns a reusable set of neural basis functions that spans the control policy space, enabling efficient zero-shot adaptation to new tasks through either projection from data or direct mapping from problem specifications. The key idea is an offline-online decomposition: basis functions are learned once during offline imitation learning, while online adaptation requires only lightweight coefficient estimation. Numerical experiments across diverse dynamics, dimensions, and cost structures show our method delivers near-optimal performance with minimal overhead when generalizing across tasks, enabling semi-global feedback policies suitable for real-time deployment.
@@ II Background
where $L:\mathbb{R}^{n}\times\mathbb{R}^{m}\times[0,T]\to\mathbb{R}$ is the running cost, $G:\mathbb{R}^{n}\to\mathbb{R}$ is the terminal cost, and $\mathcal{U}$ is the space of admissible control functions, which we assume to be sufficiently regular [6, Sec. I.3, I.8-9 ]. For simplicity and without loss of generality, we use $\boldsymbol{\eta}$} to denote the conditional variable that captures task-specific variations in $\mathcal{J}$, such as target locations, terrain types, or changes in control penalties. In practice, $\boldsymbol{\eta}$} can be complex, high-dimensional or implicit. Each choice of $\boldsymbol{\eta}$} defines a distinct optimal control problem.
@@
for each specific task. Having the control $\bf u$} to depend on both the current state ${\bf x}(t)$ and time $t$ allows for control evaluation for any state-time pair, making it more flexible and particularly useful in problems where the initial state ${\bf x}_{0}$ can vary. Classical results show that under suitable regularity assumptions, existence of optimal controls in feedback form can be guaranteed; see, e.g., [6, Sec. I.5], or [11].
@@
Function encoder (FE) [9] provides a principled framework for representing and transferring tasks in Hilbert spaces by learning a finite set of neural network basis functions. Given a Hilbert space $\mathcal{H}$, the FE learns basis functions $\{\phi_{1},\phi_{2},\dots,\phi_{p}\}$, parameterized by neural networks, such that any function $h\in\mathcal{H}$ can be approximated as
$h({\bf x})\approx\sum_{j=1}^{p}c_{j}\phi_{j}({\bf x};{\boldsymbol{\theta}}_{j}),$ (4)
@@
Online computation of coefficients for a given new problem $h$ is obtained by solving a least-squares (LS) problem
${\bf c}=\operatorname*{arg\,min}_{{\bf c}\in\mathbb{R}^{p}}\|h-\sum_{j=1}^{p}c_{j}\phi_{j}\|_{\mathcal{H}}^{2}.$ (5)
@@
${\bf c}=\begin{bmatrix}\langle\phi_{1},\phi_{1}\rangle_{\mathcal{H}}&\cdots&\langle\phi_{1},\phi_{p}\rangle_{\mathcal{H}}\\ \vdots&\ddots&\vdots\\ \langle\phi_{p},\phi_{1}\rangle_{\mathcal{H}}&\cdots&\langle\phi_{p},\phi_{p}\rangle_{\mathcal{H}}\end{bmatrix}^{-1}\begin{bmatrix}\langle h,\phi_{1}\rangle_{\mathcal{H}}\\ \vdots\\ \langle h,\phi_{p}\rangle_{\mathcal{H}}\end{bmatrix}$, (6)
@@
Let $K\subset\mathbb{R}^{n}$ be compact and let $\mathcal{H}=\{h:K\to\mathbb{R}^{n}\mid\|h\|_{\mathcal{H}}<\infty\}$ be a separable Hilbert space. For any continuous $h\in\mathcal{H}$ and any $\epsilon>0$, there exist neural network basis functions $\{\phi_{1},\phi_{2},\dots\}$, some positive integer $P$ and coefficients ${\bf c}\in\mathbb{R}^{P}$ such that
$\Big\|h-\sum_{j=1}^{P}c_{j}\,\phi_{j}\Big\|_{\mathcal{H}}<\epsilon\,\|h\|_{\mathcal{H}}.$
@@ 7 ${\boldsymbol{\gamma}}\leftarrow{\boldsymbol{\gamma}}-\beta\,\nabla_{{\boldsymbol{\gamma}}}L$;
We note that while Theorem 1 guarantees that the bias of the policy adaptation can be made arbitrarily small, (8) is a discrete form of (5), where the inner product in $\mathcal{H}$ is approximated by finite measurements via Monte Carlo integration. Given some task, denote the minimizers of (5) and (8) by $\bf c$$^{\star}$ and $\bf c$}. They differ due to finite sample approximation of the Gram matrix and right-hand side.
@@
Assume i.i.d. sampling of $({\bf x}_{i},t_{i})_{i=1}^{M}$ from the measure defining the inner product in $\mathcal{H}$. Denote by $\Phi\in\mathbb{R}^{M\times p}$ the matrix whose $i$-th row is $\phi({\bf x}_{i},t_{i})^{\top}$, $B=\tfrac{1}{M}\Phi^{\top}\Phi$ is the sample average of the Gram matrix $B^{\star}$ with $\lambda_{\min}(B^{\star})>0$. Under uniform boundedness assumptions for each $\phi$ and target function $\bf u$}, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\|{\bf c}-{\bf c}^{\star}\|_{2}\leq\frac{\lambda_{\min}(B^{\star})\sigma_{1}+\sigma_{2}}{\lambda^{2}_{\min}(B^{\star})})\sqrt{\frac{\ln(p/\delta)}{M}}=\mathcal{O}\big(M^{-1/2}\big),$
where $\sigma$ depends on the learned basis functions and $\bf u$}.
The result follows the concentration bounds for finite sample covariance matrices [14], the details of which we provide in the appendix. Together with Theorem 1, this implies asymptotic convergence of policy prediction as the number of online samples increase, validating our proposed approach for both offline training and online inference steps. Empirical validation on function encoder is also available at [9].
@@ IV RELATED WORK
ML amortizes solve time into an offline phase to produce feedback policies evaluable online—bridging local trajectory optimization and global HJB solvers. Notably, these learned policies are usually semi-global: by opting for sampling instead of meshing, they provide reliable feedback over broad regions of the state–time domain, making them scalable, though they remain task-specific in their standard form. Data-driven approaches such as [26, 27, 28, 29, 30] train policies on solver-generated trajectories and can achieve fast convergence and high accuracy. Model-based and physics-informed approaches [31, 32, 33, 34, 35, 36, 37, 38, 39, 40] leverage OC structure (objective terms, HJB residuals, Koopman representation, etc.) to provide robustness and reduce data requirements in training, though the resulting optimization can be considerably more challenging. Model-free RL [41, 42] learns feedback without an explicit mathematical model but as a trade-off is often sample-inefficient and may underperform model-based approaches [43]. Further integrations of deep learning and optimal control have also yielded promising directions such as [44, 45, 46, 47, 48], however these discussions are beyond the scope of this paper. While ML-based approaches have seen extraordinary success in solving optimal control problems, repurposing a trained model to new tasks efficiently and accurately remains an open challenge. This work aims to address that gap.
@@ V Numerical Experiments
Our FE implementation closely follows [9]. We use multi-head, multi-layer perceptrons (MLPs) for basis function parameterization. Exact training details vary for each example. Default regularization is set to $0.001$. For data preparation and ground truth comparison we rely on discretization-based open-loop solvers using SciPy or CasADi [18]. Our code is implemented in PyTorch [49] and can be found here. All experiments were conducted on an NVIDIA A100 GPU.
@@
Evaluation True Objective Loss Predicted Objective Loss Inference Method
@@ V-B Quadcopter Path Planning with Different Targets
Evaluation True Objective Loss Predicted Objective Loss Inference Method
@@ V-C Bicycle Control Under Different Obstacle Configurations
TABLE III: Numerical results for the obstacle avoidance example. We display the prediction and ground truth for the cumulative control and obstacle cost, as well as terminal state deviation measured in both $x$ and $y$ direction, all results are evaluated and averaged over settings unseen in training.
@@
We illustrate the numerical results in Table III. Note that the learned model transfers accurately to new problem settings using limited data for inference. In particular, the controller can learn to avoid the obstacle for arbitrary placement and reach the target state with high precision, demonstrating the capability of our proposed approach. In Figure 5, we display the results for $3$ worst performing subproblems out of all tested cases. Notice that even in the worst performing cases, the learned model can provide reasonable and accurate guidance to the system, as demonstrated in the close match between the predicted and ground truth trajectories.
@@ VI Conclusions
APPENDIX
Proof of Theorem 2.  We define the population LS coefficients
${\bf c}_{*}=\begin{bmatrix}\langle\phi_{1},\phi_{1}\rangle_{\mathcal{H}}&\cdots&\langle\phi_{1},\phi_{p}\rangle_{\mathcal{H}}\\ \vdots&\ddots&\vdots\\ \langle\phi_{p},\phi_{1}\rangle_{\mathcal{H}}&\cdots&\langle\phi_{p},\phi_{p}\rangle_{\mathcal{H}}\end{bmatrix}^{-1}\begin{bmatrix}\langle{\bf u},\phi_{1}\rangle_{\mathcal{H}}\\ \vdots\\ \langle{\bf u},\phi_{p}\rangle_{\mathcal{H}}\end{bmatrix}=\mathbb{E}[\Phi^{\top}\Phi]^{-1}\mathbb{E}[\Phi^{\top}{\bf u}]={\bf B}_{*}^{-1}{\bf r}_{*},$
and the corresponding sample average approximation
${\bf c}=\begin{bmatrix}\langle\phi_{1},\phi_{1}\rangle_{M}&\cdots&\langle\phi_{1},\phi_{p}\rangle_{M}\\ \vdots&\ddots&\vdots\\ \langle\phi_{p},\phi_{1}\rangle_{M}&\cdots&\langle\phi_{p},\phi_{p}\rangle_{M}\end{bmatrix}^{-1}\begin{bmatrix}\langle{\bf u},\phi_{1}\rangle_{M}\\ \vdots\\ \langle{\bf u},\phi_{p}\rangle_{M}\end{bmatrix}=[\frac{1}{M}\sum_{i=1}^{M}\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})]^{-1}[\frac{1}{M}\sum_{i=1}^{M}\Phi^{\top}({\bf x}_{i}){\bf u}({\bf x}_{i})]={\bf B}_{M}^{-1}{\bf r}_{M},$
where $\langle\cdot,\cdot\rangle_{M}$ is the Monte-Carlo approximation of the inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ using $M$ samples. We consider the following standard assumptions
  1. 1.
    $\|\Phi({\bf x})\|_{2}^{2}\leq C_{1}$ for some constant $C_{1}$ and for all $\bf x$}.
  2. 2.
    $\|\Phi^{\top}({\bf x}){\bf u}({\bf x})\|_{2}\leq C_{2}$, for some constant $C_{2}$ and for all $\bf x$}.
Note that we have
$\|{\bf B}_{*}\|_{2}\leq\|\mathbb{E}(\Phi^{\top}\Phi)\|_{2}\leq\mathbb{E}\left(\|\Phi^{\top}\Phi\|_{2}\right)=\mathbb{E}\left(\|\Phi\|_{2}^{2}\right)\leq\mathbb{E}(C_{1})=C_{1},$ (11)
where we used Jensen’s inequality in the second step, and assumption 1 in the second last step. Similarly, we have
$\|{\bf r}_{*}\|_{2}=\|\mathbb{E}(\Phi^{\top}{\bf u})\|_{2}\leq\mathbb{E}\left(\|\Phi^{\top}{\bf u}\|_{2}\right)\leq\mathbb{E}(C_{2})=C_{2},$ (12)
where we used Jensen’s inequality in the second step, and assumption 2 in the second last step.
Define random matrices
${\bf Z}_{i}=\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})-{\bf B}_{*},\quad\text{for}\ i=1,2,\dots,M.$
Then $\{{\bf Z}_{i}\}_{i=1}^{M}$ are independent, mean-zero random matrices and
$\|{\bf Z}_{i}\|_{2}\leq\|\Phi({\bf x}_{i})^{\top}\Phi({\bf x}_{i})\|_{2}+\|{\bf B}_{*}\|_{2}=\|\Phi({\bf x}_{i})\|^{2}_{2}+\|{\bf B}_{*}\|_{2}\leq 2C_{1},$
where in the first step we used triangular inequality, and in the last step we used assumption 1 and Equation 11. The variance proxy follows
$\|\mathbb{E}({\bf Z}_{i}^{2})\|_{2}$ $=\|\mathbb{E}(\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})-\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i}){\bf B}_{*}-{\bf B}_{*}\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})+{\bf B}_{*}^{2})\|_{2}$ (13)
$=\|\mathbb{E}(\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i}))-\mathbb{E}(\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})){\bf B}_{*}-{\bf B}_{*}\mathbb{E}(\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i}))+{\bf B}_{*}^{2}\|_{2}$ (14)
$=\|\mathbb{E}(\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i}))-{\bf B}_{*}^{2}\|_{2}$ (15)
$\leq\|\mathbb{E}(\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i}))^{2}\|_{2}+\|{\bf B}_{*}^{2}\|_{2}$ (16)
$\leq\mathbb{E}\|(\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i}))^{2}\|_{2}+\|{\bf B}_{*}^{2}\|_{2}$ (17)
by Jensen’s inequality,
$=\mathbb{E}\|\Phi^{\top}({\bf x}_{i})\Phi({\bf x}_{i})\|_{2}^{2}+\|{\bf B}_{*}^{2}\|_{2}$ (18)
$=\mathbb{E}\|\Phi({\bf x}_{i})\|_{2}^{4}+\|{\bf B}_{*}\|_{2}^{2}$ (19)
$\leq 2C_{1}^{2},$ (20)
where in the last step we used assumption 1 and Equation 11. Now applying vector/matrix Bernstein inequalities [53][Theorem 6.6.1], we can get the following bounds with probability $1-\delta$ for $\forall\delta\in(0,1)$,
$\|{\bf B}_{M}-{\bf B}_{*}\|_{2}\leq\sqrt{\frac{4C_{1}^{2}\ln(2p/\delta)}{M}}+\frac{4C_{1}\ln(2p/\delta)}{3M},$ (21)
where $p$ is the number of basis functions. Following similar steps and we have for the right hand side
$\|{\bf r}_{M}-{\bf r}_{*}\|_{2}\leq\sqrt{\frac{4C_{2}^{2}\ln((p+1)/\delta)}{M}}+\frac{4C_{2}\ln((p+1)/\delta)}{3M}.$ (22)
We then consider the bound for $\|{\bf B}_{M}^{-1}-{\bf B}_{*}^{-1}\|_{2}$. Assume sample size $M$ is selected large enough such that
$\|{\bf B}_{M}-{\bf B}_{*}\|_{2}\leq\frac{1}{2}\lambda_{\rm min}({\bf B}_{*}).$ (23)
Note that ${\bf B}_{M}^{-1}-{\bf B}_{*}^{-1}={\bf B}_{M}^{-1}({\bf B}_{*}-{\bf B}_{M}){\bf B}_{*}^{-1}$, we have
$\|{\bf B}_{M}^{-1}-{\bf B}_{*}^{-1}\|_{2}\leq\|{\bf B}_{M}^{-1}\|_{2}\cdot\|{\bf B}_{*}-{\bf B}_{M}\|_{2}\cdot\|{\bf B}_{*}^{-1}\|_{2}.$
By Weyl inequality,
$\lambda_{\rm min}({\bf B}_{M})\geq\lambda_{\rm min}({\bf B}_{*})+\lambda_{\rm min}({\bf B}_{M}-{\bf B}_{*}),$ (24)
also we have
$\lambda_{\rm min}({\bf B}_{M}-{\bf B}_{*})\geq-\|{\bf B}_{M}-{\bf B}_{*}\|_{2}.$ (25)
By (24) and (25), we get
$\|{\bf B}_{M}^{-1}\|_{2}$ $=\frac{1}{\lambda_{\rm min}({\bf B}_{M})}\leq\frac{1}{\lambda_{\rm min}({\bf B}_{*})+\lambda_{\rm min}({\bf B}_{M}-{\bf B}_{*})}\leq\frac{1}{\lambda_{\rm min}({\bf B}_{*})-\|{\bf B}_{M}-{\bf B}_{*}\|_{2}}$ (26)
$\leq\frac{1}{\lambda_{\rm min}({\bf B}_{*})-\frac{1}{2}\lambda_{\rm min}({\bf B}_{*})}=\frac{2}{\lambda_{\rm min}({\bf B}_{*})},$ (27)
where in the second last step, we used (23). Continuing evaluating $\|{\bf B}_{M}^{-1}-{\bf B}_{*}^{-1}\|_{2}$, and using (21), (26), and that $\|{\bf B}_{*}^{-1}\|_{2}=1/\lambda_{\rm min}({\bf B}_{*})$, we have
$\|{\bf B}_{M}^{-1}-{\bf B}_{*}^{-1}\|_{2}$ $\leq\|{\bf B}_{M}^{-1}\|_{2}\cdot\|{\bf B}_{*}-{\bf B}_{M}\|_{2}\cdot\|{\bf B}_{*}^{-1}\|_{2}$ (28)
$\leq\frac{2}{\lambda_{\rm min}({\bf B}_{*})}\left(\sqrt{\frac{4C_{1}^{2}\ln(2p/\delta)}{M}}+\frac{4C_{1}\ln(2p/\delta)}{3M}\right)\frac{1}{\lambda_{\rm min}({\bf B}_{*})}$ (29)
$=\frac{2}{\lambda_{\rm min}^{2}({\bf B}_{*})}\left(\sqrt{\frac{4C_{1}^{2}\ln(2p/\delta)}{M}}+\frac{4C_{1}\ln(2p/\delta)}{3M}\right).$ (30)
In order to bound $\|{\bf c}-{\bf c}_{*}\|_{2}$, we use the fact
$\|{\bf c}-{\bf c}_{*}\|_{2}$ $=\|{\bf B}_{M}^{-1}{\bf r}_{M}-{\bf B}_{*}^{-1}{\bf r}_{*}\|_{2}$ (31)
$=\|{\bf B}_{M}^{-1}{\bf r}_{M}-{\bf B}_{M}^{-1}{\bf r}_{*}+{\bf B}_{M}^{-1}{\bf r}_{*}-{\bf B}_{*}^{-1}{\bf r}_{*}\|_{2}$ (32)
$\leq\|{\bf B}_{M}^{-1}{\bf r}_{M}-{\bf B}_{M}^{-1}{\bf r}_{*}\|_{2}+\|{\bf B}_{M}^{-1}{\bf r}_{*}-{\bf B}_{*}^{-1}{\bf r}_{*}\|_{2}$ (33)
$\leq\|{\bf B}_{M}^{-1}\|_{2}\|{\bf r}_{M}-{\bf r}_{*}\|_{2}+\|{\bf B}_{M}^{-1}-{\bf B}_{*}^{-1}\|_{2}\|{\bf r}_{*}\|_{2},$ (34)
here, using (26), (22), (28) and (12), we obtain
$\leq\frac{2}{\lambda_{\rm min}({\bf B}_{*})}\left(\sqrt{\frac{4C_{2}^{2}\ln((p+1)/\delta)}{M}}+\frac{4C_{2}\ln((p+1)/\delta)}{3M}\right)$ (36)
$+\frac{2}{\lambda_{\rm min}^{2}({\bf B}_{*})}\left(\sqrt{\frac{4C_{1}^{2}\ln(2p/\delta)}{M}}+\frac{4C_{1}\ln(2p/\delta)}{3M}\right)C_{2}$ (37)
$=\mathcal{O}(\frac{1}{\sqrt{M}}).$ (38)